A student takes a 10 question multiple choice test (each with four options, A-D) without studying or attending class. He plans to guess. A score of 6 out of 10 is needed to pass, 8 out of 10 for high pass.
a. What are his chances of passing by guessing?
b. What are his chances of a high pass just by guessing?
c. Should he manage to pass, how likely would it be a high pass?

Answer by stanbon(48568)

A student takes a 10 question multiple choice test (each with four options, A-D) without studying or attending class. He plans to guess.
---------------------------
A score of 6 out of 10 is needed to pass,
8 out of 10 for high pass.
---------------
It is a binomial problem with p = 1/4, n=10
a. chances of passing by guessing? P(x>=6) = 1-binomcdf(10,0.25,5)=0.01928
b. chances of a high pass just by guessing?= 1-binomcdf(10,0.25,7)=0.0004158..
c. Should he manage to pass, how likely would it be a high pass?
P(high pass| pass) = P(high pass AND Pass)/P(pass)=0.0004158/0.01928=0.02157
==================
Cheers,
Stan H.